# The ghost-gluon vertex in the presence of the Gribov horizon: general kinematics

### v1

@@ Abstract
Correlation functions are important probes for the behavior of quantum field theories. Already at tree-level, the Refined Gribov Zwanziger (RGZ) effective action for Yang-Mills theories provides a good approximation for the gluon propagator, as compared to that calculated by nonperturbative methods such as Lattice Field Theory and Dyson-Schwinger Equations. However, the study of higher correlation functions of the RGZ theory is still at its beginning. In this work we evaluate the ghost-antighost-gluon vertex function in Landau gauge at one-loop level, in $d=4$ space-time dimensions for the gauge groups SU(2) and SU(3). More precisely, we extend the analysis conducted in Mintz et al. (2018) for the soft-gluon limit to an arbitrary kinematic configuration. We introduce renormalization group effects by means of a toy model for the running coupling and investigate the impact of such a model in the ultraviolet tails of our results. We find that RGZ results match fairly closely those from lattice simulations, Schwinger-Dyson equations and the Curci-Ferrari model for three different kinematic configurations. This is compatible with RGZ being a feasible theory for the strong interaction in the infrared regime.
@@ I Introduction
Since the proposal of Quantum Chromodynamics (QCD) as the theory of Strong Interactions, a long path was constructed to connect the fundamental degrees of freedom – quarks and gluons – to the observed physical states and processes. At high energies, asymptotic freedom allows for a perturbative approach that, supplemented by essential nonperturbative information in Particle Distribution Functions and fragmentation phenomena, agrees with a plethora of experimental output from high-energy particle colliders. The infrared (IR) regime however is much less amenable. Monte Carlo simulations that solve the Euclidean version of QCD on a discretized space-time lattice have by now and with great effort established that this non-Abelian theory can quantitatively describe several hadronic observables Fodor and Hoelbling (2012); Durr et al. (2008); Borsanyi et al. (2021, 2015). Nevertheless, the mechanism of color confinement is still an open question, calling for the development of continuum approaches and (semi)analytical descriptions of the infrared behavior of Strong Interactions. Among the well-developed continuum methods that try to tackle this non-perturbative regime, Schwinger-Dyson equations Bashir et al. (2012); Aguilar et al. (2016, 2023, 2022, 2021a, 2021b); Huber et al. (2021); Aguilar et al. (2021a); Fischer and Huber (2020); Huber (2020); Aguilar et al. (2019) stand out in different hadronic applications, but also the Functional Renormalization Group Berges et al. (2002); Pawlowski (2007); Corell et al. (2018); Cyrol et al. (2016); Fischer et al. (2009); Dupuis et al. (2021) and effective models Nambu and Jona-Lasinio (1961); Klevansky (1992); Gell-Mann and Levy (1960); Fukushima (2004); Schaefer et al. (2007) have been employed with partial success. Other approaches, such as the Curci-Ferrari (CF) model in Landau gauge Tissier and Wschebor (2010, 2011); Pelaez et al. (2013); Peláez et al. (2014, 2015); Gracey et al. (2019); Peláez et al. (2021); Figueroa and Peláez (2022); Barrios et al. (2020, 2021, 2022) and the screened massive expansion Siringo (2023); Siringo and Comitini (2023); Comitini et al. (2021); Comitini and Siringo (2020); Siringo and Comitini (2018) have successfully described some aspects of the IR of Yang-Mills theory by employing perturbation theory.
@@
Here we adopt another continuum approach to the nonperturbative regime of Yang-Mills theories: the Refined Gribov-Zwanziger (RGZ) theory Dudal et al. (2008, 2010). This framework, as the other continuum methods, adopts a gauge-fixed setup and is formulated from first-principles as a gauge path integral modified in the infrared by the existence of Gribov gauge copies. This idea follows the seminal work by Gribov himself Gribov (1978) and the development of local actions attained by Zwanziger Zwanziger (1989) and complemented by the emergence of dimension two condensates Dudal et al. (2008, 2010, 2019). For comprehensive reviews, the reader is referred to Vandersickel and Zwanziger (2012); Vandersickel (2011); Sobreiro and Sorella (2005) and references therein.
@@
The presence of this nonperturbative background stemming from the Gribov horizon and the condensates seems to carry plenty of information from the interacting theory, so that the remaining interaction corrections might be supposed to be small, i.e. perturbative. Even at tree-level, the RGZ gluon propagator is compatible with the deep IR behavior observed on Landau-gauge Lattice QCD data Dudal et al. (2008), while reducing to pure Yang-Mills at large energies Capri et al. (2016a) in a fully self-consistent formulation that would include solving the Gribov gap equation and the effective potential for the different condensates. Gauge-parameter independence of gauge-invariant quantities in this setup is under control within linear covariant gauges through a nilpotent Becchi-Rouet-Stora–Tyutin symmetry, and the predictions for the gluon propagator in general covariant gauges are also compatible with available lattice results Capri et al. (2015, 2016b, 2016c, 2017).
@@
In this paper, the aim is to take a step forward in the direction of establishing predictions of the RGZ theory by including radiative corrections and confronting them with lattice data and other nonperturbative approaches. This is crucial for further understanding whether this theory can be a consistent model of infrared Yang-Mills and eventually QCD. Previous results on the vertices of the (R)GZ action in the Landau gauge may be found in Ref. Gracey (2012). Here, in particular, we compute the general kinematics of the gluon-ghost-antighost vertex at one loop order within the RGZ framework, extending previous work on the soft gluon limit of the same correlation function Mintz et al. (2018). The present study enables the investigation of the role played by auxiliary fields in quantitative results in RGZ. In this specific correlation function, we show that the contribution of auxiliary fields, in the form of new vertices and mixed propagators, represents a smaller effect – less than $10\%$ percent – as compared to that of the presence of nonzero pole masses for the gluon. Moreover, we include running effects via an infrared model, with a freezing coupling constant. Even though this is not self-consistently obtained within RGZ, this may be a reasonable approximation, because of the presence of the massive parameters from the horizon and the condensates. Overall we show that the RGZ results are fully compatible with available lattice data for SU(2) and SU(3), as well as other continuum approaches, such as the infrared-safe Curci-Ferrari model in Landau gauge Pelaez et al. (2013); Barrios et al. (2020) and the Dyson-Schwinger equations Aguilar et al. (2019).
@@
This paper is organized as follows. In Sec. II, the formalism of the refined Gribov-Zwanziger framework is presented. In Sec. III, the ghost-antighost-gluon vertex is defined and its general structure in the RGZ theory is discussed. Our computation of generic momentum configurations of the one-loop ghost-gluon vertex is presented in Sec. IV, including the IR running coupling model adopted. Sec. V collects our results along with the corresponding analysis and comparisons with alternative approaches. In Sec. VI we discuss the potential influence of the IR running coupling model on the ultraviolet (UV) tails of our results. We have included perspectives and final remarks in Section VII.
@@ II The Refined Gribov-Zwanziger action in Landau gauge
The Gribov-Zwanziger theory is a framework for making sense of the gauge fixed Yang-Mills theory in the non-perturbative regime. It grew from Gribov’s observation that gauge fixing in the Landau Gauge fails in the strongly coupled regime Gribov (1978); Zwanziger (1989) (see also Vandersickel (2011); Vandersickel and Zwanziger (2012); Sobreiro and Sorella (2005) for reviews), in the sense that gauge copies still exist after the fixing. These are called Gribov copies and they appear in fact in any covariant gauge. The solution proposed by Gribov was to restrict the integration region of the gauge field path integral to the neighborhood containing the perturbative vacuum and bounded by the field configurations associated with the first Gribov copies, called the Gribov Horizon. In the Landau gauge $\partial_{\mu}A^{a}_{\mu}=0$ the Gribov horizon is defined by the set of fields $A^{a}_{\mu}$ for which the equation ${\cal M}^{ab}(A)\alpha^{b}=0$ has a solution, where
@@
Zwanziger Zwanziger (1989) was able to recast Gribov’s construction in the form of an effective theory whose restriction to the Gribov horizon is implemented at the level of the action. The resulting action is known as the nonlocal Gribov-Zwanziger action and defined in the Landau gauge in dimension $D$ by
$S_{nlGZ}=\int d^{D}x\;\left(\frac{1}{4}F^{a}_{\mu\nu}F^{a}_{\mu\nu}+ib^{a}\partial_{\mu}A^{a}_{\mu}+\bar{c}^{a}{\cal M}(A)^{ab}c^{b}\right)+\gamma^{D}H(A)-\gamma^{D}VD\left(N^{2}-1\right)$ (2)
@@
The vacuum defined by the gap equation is unstable and favors the formation of new condensates Dudal et al. (2008) that are related to mass scales of the gauge fields $\langle A^{2}\rangle\sim m^{2}$ and auxiliary fields $\langle\bar{\phi}\phi\rangle\sim\mu^{2}$. The incorporation of these condensates in the effective action formulation leads to a modified theory known as the Refined Gribov-Zwanziger theory, defined by
@@
It is worth pointing out the remarkable fact discovered in Capri et al. (2016c) that it is possible to recast this theory making it suitable for any linearly covariant gauge, so that the resulting RGZ theory can be made BRST invariant. This is done by replacing the gauge field $A$ by a gauge invariant composite field $A^{h}$ Lavelle and McMullan (1997); Capri et al. (2016c)
@@
where $\xi^{a}$ is the Stueckelberg field discussed in Lavelle and McMullan (1997); Capri et al. (2016d, b, 2015, c, 2017). One also needs to impose a transversality constraint on $A^{h}$,
@@ III The three-point ghost-gluon correlation function
Let us now start discussing the ghost-gluon correlation function and establish some notation. We first make some general remarks about the connected ghost-antighost-gluon three-point function in the RGZ theory and how the mixed propagators affect its building blocks. Next, for completeness, we quote the results obtained in Mintz et al. (2018) for the soft-gluon limit. Our original results for the ghost-antighost-gluon correlation function for arbitrary momenta are left for the next section.
@@ III.1 General structure of the $A\bar{c}c$ vertex
Figure 1: Feynman diagram expansion up to one-loop order for the ghost-gluon vertex in the Refined Gribov-Zwanziger. Dashed lines represent ghosts and antighosts, while the curly lines stand for gluons. Full simple and double lines, that only appear in mixed propagators, correspond to the auxiliary fields $\varphi$ and $\bar{\varphi}$, respectively. The roman numbers identifying the one-loop diagrams will be used as reference in the results section and in the appendices. The diagrams were generated by means of the Jaxodraw interface Binosi et al. (2009).
@@
the one-loop connected function (14) can then be decomposed as
$\left\langle{A_{\nu}^{a}(k)\,\bar{c}^{b}(p)\,c^{c}(q)}\right\rangle$ $=G(p)G(q)D_{AA}(k)P^{\perp}_{\mu\nu}(k)\left\{\Gamma_{A^{a}_{\mu}\bar{c}^{b}c^{c}}(k,p,r)+\frac{2g\gamma^{2}f^{ade}}{k^{2}+M^{2}}\Gamma_{\bar{c}^{b}c^{c}\varphi^{de}_{\mu}}(k,p,r)\right\}_{q=-p-k},$ (16)
@@
$\Gamma_{A^{a}_{\nu}\bar{c}^{b}c^{c}}(k,p,r):=\frac{\delta^{3}\Gamma}{\delta A_{\nu}^{a}(-k)\delta\bar{c}^{b}(-p)\delta c^{c}(-q)}$ (18)
@@
$\Gamma_{\bar{c}^{b}c^{c}\varphi^{de}_{\mu}}(k,p,r):=\frac{\delta^{3}\Gamma}{\delta\bar{c}^{b}(-p)\delta c^{c}(-q)\delta\varphi_{\nu}^{de}(-k)}.$ (19)
Note the presence of the $\Gamma_{\bar{c}c\varphi}$ and $\Gamma_{\bar{c}c\bar{\varphi}}$ vertex functions in the RGZ theory. They are zero at tree-level. In the case of $\Gamma_{\bar{c}c\varphi}$ its first nontrivial terms are given by diagrams (IV.1) and (IV.2) of fig. 1. Notice that, since $\Gamma_{\bar{c}c\varphi}=\Gamma_{\bar{c}c\bar{\varphi}}$, in fig. 1 we have included the diagrams associated with the former quantity only. The contribution of $\Gamma_{\bar{c}c\bar{\varphi}}$ to $\left\langle{A_{\nu}^{a}(k)\,\bar{c}^{b}(p)\,c^{c}(q)}\right\rangle$ is introduced by multiplying the contribution of $\Gamma_{\bar{c}c\varphi}$ by a factor of two, as can be observed in eq. 16.
@@
where we adopt the same notation as in Aguilar et al. (2019) and consider all momenta as incoming. At tree-level, the scalar structure functions are $B_{1}(k,p)=1$ and $B_{2}(k,p)=0$.
@@
$f^{ade}\Gamma_{\bar{c}^{b}c^{c}\varphi^{de}_{\mu}}(k,p,r)=-igf^{abc}\left(p_{\mu}C_{1}(k,p)+k_{\mu}C_{2}(k,p)\right)-igd^{abc}\left(p_{\mu}C^{\prime}_{1}(k,p)+k_{\mu}C^{\prime}_{2}(k,p)\right)$ (21)
@@
In order to make contact with results from Monte Carlo lattice simulations, let us consider the scalar quantity Cucchieri et al. (2006)
@@
There are clearly some differences between perturbative YM and RGZ calculations of the vertex function. The first of them is the modification of the gluon propagator brought about by the restriction to the Gribov horizon, which can be understood as the appearance of a pair of generally complex conjugate poles. A second difference is the presence of the tree-level $A\bar{\varphi}\varphi$ vertex, which couples the gluon to the auxiliary Zwanziger fields. This allows not only diagrams with auxiliary fields running in the internal loops, but also in the external legs, as long as the external propagator is a mixed one like, for example, $\left\langle{A\varphi}\right\rangle$. This possibility is realized in (16), giving rise to the contributions $\Gamma_{\varphi\bar{c}c}$ and $\Gamma_{\bar{\varphi}\bar{c}c}$, not present in perturbative YM theory. Furthermore, note that these mixed contributions only appear from one-loop order onwards, as such vertices are absent from the classical action (II). Being finite, they do not spoil the stability of the action, in agreement with the Quantum Action Principle Piguet and Sorella (1995). Finally, we note that the presence of such vertices involving the auxiliary Zwanziger fields can be thought as effective momentum-dependent gluonic interaction terms. This claim can be explicitly shown to be true, starting from the tree level action, by considering the nonlocal version of the RGZ theory (whose action can be obtained from (II) by integrating out the Zwanziger auxiliary fields) and expanding the inverse of the Faddeev-Popov operator in powers of the coupling. In this regard, we note that it is reasonable to consider the RGZ effective action not simply as an effective model with propagators given by a ratio of polynomials, but rather a renormalizable effective theory with nontrivial gluonic propagators and self-interactions, which are related to each other.

### v2

@@ Abstract
Correlation functions are important probes for the behavior of quantum field theories. Already at tree-level, the Refined Gribov Zwanziger (RGZ) effective action for Yang-Mills theories provides a good approximation for the gluon propagator, as compared to that calculated by nonperturbative methods such as Lattice Field Theory and Dyson-Schwinger Equations. However, the study of higher correlation functions of the RGZ theory is still at its beginning. In this work we evaluate the ghost-gluon vertex function in Landau gauge at one-loop level, in $d=4$ space-time dimensions for the gauge groups SU(2) and SU(3). More precisely, we extend the analysis conducted in [1] for the soft-gluon limit to an arbitrary kinematic configuration. We introduce renormalization group effects by means of a toy model for the running coupling and investigate the impact of such a model in the ultraviolet tails of our results. We find that RGZ results match fairly closely those from lattice simulations, Schwinger-Dyson equations and the Curci-Ferrari model for three different kinematic configurations. This is compatible with RGZ being a feasible theory for the strong interaction in the infrared regime.
@@ I Introduction
Since the proposal of Quantum Chromodynamics (QCD) as the theory of Strong Interactions, a long path was constructed to connect the fundamental degrees of freedom – quarks and gluons – to the observed physical states and processes. At high energies, asymptotic freedom allows for a perturbative approach that, supplemented by essential nonperturbative information in Particle Distribution Functions and fragmentation phenomena, agrees with a plethora of experimental output from high-energy particle colliders. The infrared (IR) regime however is much less amenable. Monte Carlo simulations that solve the Euclidean version of QCD on a discretized space-time lattice have by now and with great effort established that this non-Abelian theory can quantitatively describe several hadronic observables [2, 3, 4, 5]. Nevertheless, the mechanism of color confinement is still an open question, calling for the development of continuum approaches and (semi)analytical descriptions of the infrared behavior of Strong Interactions. Among the well-developed continuum methods that try to tackle this non-perturbative regime, Schwinger-Dyson equations [6, 7, 8, 9, 10, 11, 12, 10, 13, 14, 15] stand out in different hadronic applications, but also the Functional Renormalization Group [16, 17, 18, 19, 20, 21] and effective models [22, 23, 24, 25, 26] have been employed with partial success. Other approaches, such as the Curci-Ferrari (CF) model in Landau gauge [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] and the screened massive expansion [38, 39, 40, 41, 42] have successfully described some aspects of the IR of Yang-Mills theory by employing perturbation theory.
@@
Here we adopt another continuum approach to the nonperturbative regime of Yang-Mills theories: the Refined Gribov-Zwanziger (RGZ) theory [43, 44]. This framework, as the other continuum methods, adopts a gauge-fixed setup and is formulated from first-principles as a gauge path integral modified in the infrared by the existence of Gribov gauge copies. This idea follows the seminal work by Gribov himself [45] and the development of local actions attained by Zwanziger [46] and complemented by the emergence of dimension two condensates [43, 44, 47]. For comprehensive reviews, the reader is referred to [48, 49, 50] and references therein.
@@
The presence of this nonperturbative background stemming from the Gribov horizon and the condensates seems to carry plenty of information from the interacting theory, so that the remaining interaction corrections might be supposed to be small, i.e. perturbative. Even at tree-level (but also at 1-loop level, as discussed in [51]), the RGZ gluon propagator is compatible with the deep IR behavior observed on Landau-gauge Lattice QCD data [43], while reducing to pure Yang-Mills at large energies [52] in a fully self-consistent formulation that would include solving the Gribov gap equation and the effective potential for the different condensates. Gauge-parameter independence of gauge-invariant quantities in this setup is under control within linear covariant gauges through a nilpotent Becchi-Rouet-Stora–Tyutin symmetry, and the predictions for the gluon propagator in general covariant gauges are also compatible with available lattice results [53, 54, 55, 56].
@@
In this paper, the aim is to take a step forward in the direction of establishing predictions of the RGZ theory by including radiative corrections and confronting them with lattice data and other nonperturbative approaches. This is crucial for further understanding whether this theory can be a consistent model of infrared Yang-Mills and eventually QCD. One-loop corrections to the RGZ gluon propagator have been recently reported in [51], whilw a corrected propagator for a scalar field coupled to the RGZ action has been studied in [57]. Previous results on the vertices of the (R)GZ action in the Landau gauge may be found in Ref. [58]. Here, in particular, we compute the general kinematics of the ghost-gluon vertex at one loop order within the RGZ framework, extending previous work on the soft gluon limit of the same correlation function [1]. The present study enables the investigation of the role played by auxiliary fields in quantitative results in RGZ. In this specific correlation function, we show that the contribution of auxiliary fields, in the form of new vertices and mixed propagators, represents a smaller effect – less than $10\%$ percent – as compared to that of the presence of nonzero pole masses for the gluon. Moreover, we include running effects via an infrared model, with a freezing coupling constant. Even though this is not self-consistently obtained within RGZ, this may be a reasonable approximation, because of the presence of the massive parameters from the horizon and the condensates. Overall we show that the RGZ results are fully compatible with available lattice data for SU(2) and SU(3), as well as other continuum approaches, such as the infrared-safe Curci-Ferrari model in Landau gauge [29, 35] and the Dyson-Schwinger equations [15].
@@
This paper is organized as follows. In Sec. II, the formalism of the refined Gribov-Zwanziger framework is presented. In Sec. III, the ghost-gluon vertex is defined and its general structure in the RGZ theory is discussed. Our computation of generic momentum configurations of the one-loop ghost-gluon vertex is presented in Sec. IV, including the IR running coupling model adopted. Sec. V collects our results along with the corresponding analysis and comparisons with alternative approaches. In Sec. VI we discuss the potential influence of the IR running coupling model on the ultraviolet (UV) tails of our results. We have included perspectives and final remarks in Section VII.
@@ II The Refined Gribov-Zwanziger action in Landau gauge
The Gribov-Zwanziger theory is a framework for making sense of the gauge fixed Yang-Mills theory in the non-perturbative regime. It grew from Gribov’s observation that gauge fixing in the Landau Gauge fails in the strongly coupled regime [45, 46] (see also [49, 48, 50] for reviews), in the sense that gauge copies still exist after the fixing. These are called Gribov copies and they appear in fact in any covariant gauge. The solution proposed by Gribov was to restrict the integration region of the gauge field path integral to the neighborhood containing the perturbative vacuum and bounded by the field configurations associated with the first Gribov copies, called the Gribov Horizon. In the Landau gauge $\partial_{\mu}A^{a}_{\mu}=0$ the Gribov horizon is defined by the set of fields $A^{a}_{\mu}$ for which the equation ${\cal M}^{ab}(A)\alpha^{b}=0$ has a solution, where
@@
Zwanziger [46] was able to recast Gribov’s construction in the form of an effective theory whose restriction to the Gribov horizon is implemented at the level of the action. The resulting action is known as the nonlocal Gribov-Zwanziger action and defined in the Landau gauge in dimension $D$ by
$S_{nlGZ}=\int d^{D}x\;\left(\frac{1}{4}F^{a}_{\mu\nu}F^{a}_{\mu\nu}+ib^{a}\partial_{\mu}A^{a}_{\mu}-\bar{c}^{a}{\cal M}(A)^{ab}c^{b}\right)+\gamma^{D}H(A)-\gamma^{D}VD\left(N^{2}-1\right)$ (2)
@@
The vacuum defined by the gap equation is unstable and favors the formation of new condensates [43] that are related to mass scales of the gauge fields $\langle A^{2}\rangle\sim m^{2}$ and auxiliary fields $\langle\bar{\phi}\phi\rangle\sim\mu^{2}$. The incorporation of these condensates in the effective action formulation leads to a modified theory known as the Refined Gribov-Zwanziger theory, defined by
@@
It is worth pointing out the remarkable fact discovered in [55] that it is possible to recast this theory making it suitable for any linearly covariant gauge, so that the resulting RGZ theory can be made BRST invariant. This is done by replacing the gauge field $A$ by a gauge invariant composite field $A^{h}$ [59, 55]
@@
where $\xi^{a}$ is the Stueckelberg field discussed in [59, 60, 54, 53, 55, 56]. One also needs to impose a transversality constraint on $A^{h}$,
@@ III The three-point ghost-gluon correlation function
Let us now start discussing the ghost-gluon correlation function and establish some notation. We first make some general remarks about the connected gluon-antighost-ghost three-point function in the RGZ theory and how the mixed propagators affect its building blocks. Next, for completeness, we quote the results obtained in [1] for the soft-gluon limit. Our original results for the gluon-antighost-ghost correlation function for arbitrary momenta are left for the next section.
@@ III.1 General structure of the $A\bar{c}c$ vertex
Figure 1: Feynman diagram expansion up to one-loop order for the ghost-gluon vertex in the Refined Gribov-Zwanziger theory. Dashed lines represent ghosts and antighosts, while the curly lines stand for gluons. Full simple and double lines, that only appear in mixed propagators, correspond to the auxiliary fields $\varphi$ and $\bar{\varphi}$, respectively. Of course, vertices are evaluated at tree-level order by employing the expressions provided in appendix A. Note that as for the diagrams (IV.1) and (IV.2) the external leg involving the gluon corresponds to a mixed propagator of the form $\left\langle{A\varphi}\right\rangle$. The roman numbers identifying the one-loop diagrams will be used as reference in the results section and in the appendices. The diagrams were generated by means of the Jaxodraw interface [61].
@@
the one-loop connected function (14) can then be decomposed as [^2] This expression has already been derived in [1]. However, the result showed here differs from it by a factor of (-2i) in the coefficient of $\Gamma_{\bar{c}^{b}c^{c}\varphi_{\mu}^{de}}$ as a consequence of a different choice for the Feynman rules. Of course, this is just a matter of convention and does not impact on the evaluation of the vertex function.
$\left\langle{A_{\nu}^{a}(k)\,\bar{c}^{b}(p)\,c^{c}(q)}\right\rangle$ $=G(p)G(q)D_{AA}(k)P^{\perp}_{\mu\nu}(k)\left\{\Gamma_{A^{a}_{\mu}\bar{c}^{b}c^{c}}(k,p,r)-\frac{2ig\gamma^{2}f^{ade}}{k^{2}+M^{2}}\Gamma_{\bar{c}^{b}c^{c}\varphi^{de}_{\mu}}(k,p,r)\right\}_{q=-p-k},$ (16)
@@
$\Gamma_{A^{a}_{\mu}\bar{c}^{b}c^{c}}(k,p,r):=\frac{\delta^{3}\Gamma}{\delta A_{\mu}^{a}(-k)\delta\bar{c}^{b}(-p)\delta c^{c}(-q)}$ (18)
@@
$\Gamma_{\bar{c}^{b}c^{c}\varphi^{de}_{\mu}}(k,p,r):=\frac{\delta^{3}\Gamma}{\delta\bar{c}^{b}(-p)\delta c^{c}(-q)\delta\varphi_{\mu}^{de}(-k)}.$ (19)
Note the presence of the $\Gamma_{\bar{c}c\varphi}$ and $\Gamma_{\bar{c}c\bar{\varphi}}$ vertex functions in the RGZ theory, which are zero at tree-level order. In the case of $\Gamma_{\bar{c}c\varphi}$ its first nontrivial terms are given by diagrams (IV.1) and (IV.2) of fig. 1, where the external leg involving the gluon field corresponds to a propagator of the type $\left\langle{A\varphi}\right\rangle$ connected to a $\bar{\varphi}\varphi A$ tree level vertex. In contrast, the vertex functions $\Gamma_{\bar{c}c\omega}$ and $\Gamma_{\bar{c}c\bar{\omega}}$ do not contribute to the ghost-gluon vertex due to the absence of mixed propagators of the type $\left\langle{A\omega}\right\rangle$ within the RGZ theory. Notice also that, since $\Gamma_{\bar{c}c\varphi}=\Gamma_{\bar{c}c\bar{\varphi}}$, in fig. 1 we have included the diagrams associated with the former quantity only. The contribution of $\Gamma_{\bar{c}c\bar{\varphi}}$ to $\left\langle{A_{\nu}^{a}(k)\,\bar{c}^{b}(p)\,c^{c}(q)}\right\rangle$ is introduced by multiplying the contribution of $\Gamma_{\bar{c}c\varphi}$ by a factor of two, as can be observed in eq. 16.
@@
where we adopt the same notation as in [15] and consider all momenta as incoming. At tree-level, the scalar structure functions are $B_{1}(k,p)=1$ and $B_{2}(k,p)=0$.
@@
$f^{ade}\Gamma_{\bar{c}^{b}c^{c}\varphi^{de}_{\mu}}(k,p,r)=gf^{abc}\left(p_{\mu}C_{1}(k,p)+k_{\mu}C_{2}(k,p)\right)+gd^{abc}\left(p_{\mu}C^{\prime}_{1}(k,p)+k_{\mu}C^{\prime}_{2}(k,p)\right)$ (21)
@@
In order to make contact with results from Monte Carlo lattice simulations, let us consider the scalar quantity [62]
@@
There are clearly some differences between perturbative YM and RGZ calculations of the vertex function. The first of them is the modification of the gluon propagator brought about by the restriction to the Gribov horizon, which can be understood as the appearance of a pair of generally complex conjugate poles. A second difference is the presence of the tree-level $A\bar{\varphi}\varphi$ vertex, which couples the gluon to the auxiliary Zwanziger fields. This allows not only diagrams with auxiliary fields running in the internal loops, but also in the external legs, as long as the external propagator is a mixed one like, for example, $\left\langle{A\varphi}\right\rangle$. This possibility is realized in (16), giving rise to the contributions $\Gamma_{\varphi\bar{c}c}$ and $\Gamma_{\bar{\varphi}\bar{c}c}$, not present in perturbative YM theory. Furthermore, note that these mixed contributions only appear from one-loop order onwards, as such vertices are absent from the classical action (II). Being finite, they do not spoil the stability of the action, in agreement with the Quantum Action Principle [63]. Finally, we note that the presence of such vertices involving the auxiliary Zwanziger fields can be thought as effective momentum-dependent gluonic interaction terms. This claim can be explicitly shown to be true, starting from the tree level action, by considering the nonlocal version of the RGZ theory (whose action can be obtained from (II) by integrating out the Zwanziger auxiliary fields) and expanding the inverse of the Faddeev-Popov operator in powers of the coupling. In this regard, we note that it is reasonable to consider the RGZ effective action not simply as an effective model with propagators given by a ratio of polynomials, but rather a renormalizable effective theory with nontrivial gluonic propagators and self-interactions, which are related to each other.